# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v5

@@ 7.3. General factorization theorem
    $\Delta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\equiv\eta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\otimes\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}}(z))\quad\mod{(U\otimes U_{+})}[\negthinspace[z]\negthinspace].$
@@
where the second and fifth equalities follow from part (1) of the present theorem, and the fourth equality from Theorem 4.9 and the group-like property of the series $\boldsymbol{\psi}(z),\boldsymbol{\phi}(z)$ (see, e.g., [FR99, Lemma 1], originally proven in [Dam98]). ∎
As a consequence, we also obtain a generalization of Corollary 5.5. Let $V$ be an irreducible finite dimensional $U_{q}L\mathfrak{sl}_{2}$-module and $\eta_{\mathbf{c},\mathbf{s}}^{*}(V)$ its restriction to $\mathcal{O}_{q}^{\mathbf{c}}$ via the embedding (2.20), with $s_{0},s_{1}$ arbitrary. Let $w\in\eta_{\mathbf{c},\mathbf{s}}^{*}(V)$ be a generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector. Let $\boldsymbol{D}_{\eta_{\mathbf{c},\mathbf{s}}^{*}(V),w}(z)\in{\mathbb{C}}[\negthinspace[z]\negthinspace]$ be defined by $\boldsymbol{\grave{\Theta}}(z)\cdot w=\boldsymbol{D}_{\eta_{\mathbf{c},\mathbf{s}}^{*}(V),w}(z)w$.
@@ 8.1. Quantum affine algebras
(8.13) $[x^{+}_{i,k},x^{-}_{j,l}]=$ $\ \delta_{ij}\textstyle\frac{1}{q-q^{-1}}(\psi_{i,k+l}-\phi_{i,k+l}),$
@@
$\operatorname{deg}^{\mathrm{Dr}}x^{\pm}_{i,k}=\pm\alpha_{i},\quad\operatorname{deg}^{\mathrm{Dr}}h_{i,k}=\operatorname{deg}^{\mathrm{Dr}}\mathbb{K}_{i}=0\qquad(i\in\mathbb{I}_{0}).$
@@
(8.17) $U_{-}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\forall j\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\leq 0,\ \exists j\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y<0\rangle,$
@@
(8.20) $U_{\neq i,-}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\leq 0,\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y<0\rangle,$
(8.21) $U_{i+,0}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y>0,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y=0\rangle,$
(8.22) $U_{i-,0}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y<0,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y=0\rangle.$
Finally, let $U^{i}_{+}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ generated by $U_{\neq i,+}$ and $U_{i+,0}$. It is the span of all elements which are: either (a) of positive degree along $\alpha_{i}$ and non-negative degree along all the other simple roots, or (b) of arbitrary degree along $\alpha_{i}$ and positive $Q_{i}$-degree. Similarly, let $U^{i}_{-}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ generated by $U_{\neq i,-}$ and $U_{i-,0}$.
@@ 8.3. Generalized $q$-Onsager algebras
(8.32) $R(r_{1},r_{2}\mid s;i,j)=$ $\ q^{2}\mathfrak{c}_{i}\mathfrak{C}^{r_{1}}\big{(}-\sum_{p\geq 0}q^{2p}[2][\Theta_{i,r_{2}-r_{1}-2p-1},B_{j,l-1}]_{q^{-2}}\mathfrak{C}^{p+1}$
(8.33) $\ -\sum_{p\geq 1}q^{2p-1}[2][B_{j,l},\Theta_{i,r_{2}-r_{1}-2p}]_{q^{-2}}\mathfrak{C}^{p}-[B_{j,l},\Theta_{i,r_{2}-r_{1}}]_{q^{-2}},$
@@ 8.4. Coideal structures
giving $\mathcal{O}_{q}^{\mathbf{c}}(\widehat{{\mathfrak{g}}})$ the structure of a right coideal subalgebra of $U_{q}L\mathfrak{g}$. There is also a generic version of this embedding
@@ 9.4. Weak compatibility
Let us write $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ as a sum of $Q$-homogeneous terms. We say that a term is irrelevant if its degree is not ±$\pm\alpha_{i}$. The proposition below is our first result on the compatibility of the two braid group actions.
Proposition 9.6.
@@
for $2\leq k\leq i$. Therefore, Lemmas 8.4 and 9.3-9.4 imply that
@@
Let $\mu$ denote the composition of the east and south arrows, and $\nu$ the composition of the south and east arrows in (9.1). We have $\mu(B_{0})=\nu(B_{0})$ and, by Proposition 9.6, $\mu(B_{-1})\equiv\nu(B_{-1})$ modulo $U_{d_{i}=1,+}$. Since $B_{0}$ and $B_{-1}$ generate $\mathcal{O}_{q}$, and $\mu(B_{0})$ contains terms of degree ±$\pm\alpha_{i}$, it follows that the diagram commutes modulo $U_{\neq i,+}$. ∎
@@ 9.5. Strong compatibility
The result holds for $r=-1,0$ by Proposition 9.6. The general case is proven using upward and downward induction. We explicitly do the former, the latter being analogous. Relation (8.36) implies that
(9.34) $A_{i,r+1}=[2]_{q}^{-1}[H_{i,1},A_{i,r}]+\mathfrak{C}A_{i,r-1}.$
@@ 9.7. Factorization theorem in higher rank
We can now generalize the factorization and coproduct theorems to higher ranks.
@@
(9.70) $\Delta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\otimes\eta(\boldsymbol{\grave{\Theta}}_{i}(z))\quad\mod{(U\otimes U_{+})}[\negthinspace[z]\negthinspace].$
@@
Thanks to Corollary 9.15, Corollaries 5.1 and 5.5 carry over to higher rank, as formulated in the introduction (Theorem 1.3).

### v6

@@ 7.3. General factorization theorem
    $\Delta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\equiv\eta_{\mathbf{c},\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\otimes\eta_{\mathbf{c},0}(\boldsymbol{\grave{\Theta}}(z))\quad\mod{(\mathcal{O}_{q}^{\mathbf{c},\mathbf{s}}\otimes U_{+})}[\negthinspace[z]\negthinspace].$
@@
where the second and fifth equalities follow from part (1) of the present theorem, and the fourth equality from Theorem 4.9 and the group-like property of the series $\boldsymbol{\phi}^{+}(z),\boldsymbol{\phi}^{-}(z)$ (see, e.g., [FR99, Lemma 1], originally proven in [Dam98]). ∎
As a consequence, we also obtain a generalization of Theorem 5.5. Let $V$ be an irreducible finite dimensional $U_{q}L\mathfrak{sl}_{2}$-module and $\eta_{\mathbf{c},\mathbf{s}}^{*}(V)$ its restriction to $\mathcal{O}_{q}^{\mathbf{c}}$ via the embedding (2.21), with $s_{0},s_{1}$ arbitrary. Let $w\in\eta_{\mathbf{c},\mathbf{s}}^{*}(V)$ be a generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector. Let $\boldsymbol{D}_{\eta_{\mathbf{c},\mathbf{s}}^{*}(V),w}(z)\in{\mathbb{C}}[\negthinspace[z]\negthinspace]$ be defined by $\boldsymbol{\grave{\Theta}}(z)\cdot w=\boldsymbol{D}_{\eta_{\mathbf{c},\mathbf{s}}^{*}(V),w}(z)w$.
@@ 8. Affine QSP of split type
Our next goal is to generalize the factorization and coproduct formulae (Theorem 7.5) to QSP coideal subalgebras of higher rank. This will be achieved in §9. Below we recall the definitions of quantum affine algebras and generalized $q$-Onsager algebras, as well as certain facts about braid group actions.
@@ 8.1. Quantum affine algebras
(8.13) $[x^{+}_{i,k},x^{-}_{j,l}]=$ $\ \delta_{ij}\textstyle\frac{1}{q-q^{-1}}(\phi^{+}_{i,k+l}-\phi^{-}_{i,k+l}),$
@@
$\operatorname{deg}^{\mathrm{Dr}}x^{\pm}_{i,k}=\pm\alpha_{i},\quad\operatorname{deg}^{\mathrm{Dr}}h_{i,k}=\operatorname{deg}^{\mathrm{Dr}}\mathbb{K}_{j}=0\qquad(i\in\mathbb{I}_{0},\ j\in\mathbb{I}).$
@@
(8.19) $U_{i+,0}=$ $\ \langle y\in U_{q}^{\mathbb{K}}L\mathfrak{g}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y>0,\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y=0\rangle.$
Finally, let $U^{i}_{+}$ be the subalgebra of $U_{q}^{\mathbb{K}}L\mathfrak{g}$ generated by $U_{\neq i,+}$ and $U_{i+,0}$. It is the span of all elements which are: either (a) of positive degree along $\alpha_{i}$ and non-negative degree along all the other simple roots, or (b) of arbitrary degree along $\alpha_{i}$ and positive $Q_{i}$-degree.
@@ 8.3. Generalized $q$-Onsager algebras
(8.29) $R(r_{1},r_{2}\mid s;i,j)=$ $\ q^{2}\mathfrak{c}_{i}\mathfrak{C}^{r_{1}}\big{(}-\sum_{p\geq 0}q^{2p}[2][\Theta_{i,r_{2}-r_{1}-2p-1},A_{j,s-1}]_{q^{-2}}\mathfrak{C}^{p+1}$
(8.30) $\ -\sum_{p\geq 1}q^{2p-1}[2][A_{j,s},\Theta_{i,r_{2}-r_{1}-2p}]_{q^{-2}}\mathfrak{C}^{p}-[A_{j,s},\Theta_{i,r_{2}-r_{1}}]_{q^{-2}}\big{)},$
@@ 8.4. Coideal structures
giving $\mathcal{O}_{q}^{\mathbf{c}}(\widehat{{\mathfrak{g}}})$ the structure of a right coideal subalgebra of $U_{q}L\mathfrak{g}$. Abbreviate $\mathcal{O}_{q}^{\mathbf{c},\mathbf{s}}(\widehat{{\mathfrak{g}}})=\eta_{\mathbf{c},\mathbf{s}}(\mathcal{O}_{q}^{\mathbf{c}}(\widehat{{\mathfrak{g}}}))$. There is also a generic version of this embedding
@@ 9. Compatibility of braid group actions
Remark 9.1.
In rank 1, it was sufficient to work with $q$-Onsager algebra with fixed parameters $\mathbf{c}$ and $\mathbf{s}$. In higher rank, we first need to pass to the universal version of the generalized $q$-Onsager algebra. This is necessitated by our analysis of the braid group action, which is non-trivial on the central elements $\mathbb{K}_{i}$. However, in the main result of this section, Corollary 9.16, we can specialize the parameters again.
@@ 9.4. Weak compatibility
Let us write $\eta(\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}))$ as a sum of $Q$-homogeneous terms. We say that a term is irrelevant if its degree is not ±$\pm\alpha_{i}$. The theorem below is our first result on the compatibility of the two braid group actions. It is the essential part of Theorem 1.4 from the Introduction.
Theorem 9.7.
@@
for $2\leq k\leq i$. Therefore, Lemmas 8.4 and 9.4-9.5 imply that
@@
[…]
@@ 9.5. Strong compatibility
The result holds for $r=-1,0$ by Theorem 9.7. The general case is proven using upward and downward induction. We explicitly do the former, the latter being analogous. Relation (8.33) implies that
(9.34) $A_{i,r+1}=[2]^{-1}[H_{i,1},A_{i,r}]+\mathfrak{C}A_{i,r-1}.$
@@ 9.7. Factorization theorem in higher rank
We can now generalize the factorization and coproduct theorems to higher ranks. Let us choose parameters $\mathbf{c},\mathbf{s}$, which determines the coideal structure through (8.41).
@@
(9.70) $\Delta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\otimes\eta(\boldsymbol{\grave{\Theta}}_{i}(z))\quad\mod{(\mathcal{O}_{q}^{\mathbf{c},\mathbf{s}}(\widehat{{\mathfrak{g}}})\otimes U_{+})}[\negthinspace[z]\negthinspace].$
@@
Thanks to Corollary 9.16, Corollary 5.1 and Theorem 5.5 carry over to higher rank, as formulated in the introduction (Theorem 1.3).